OpenStudy (anonymous):

3/8=x/12 solve for x solve for x x+1/8=2x-5/2 3x/3x-36=-1/x solve for x 2x/16x+30=-1/x solve for x medals for all

OpenStudy (anonymous):

x=9/2

OpenStudy (anonymous):

x=21/8

OpenStudy (anonymous):

x=+6,-6

OpenStudy (anonymous):

what order are you going in because i have a list of answers and none of them are on there not being rude but im not understanding

OpenStudy (anonymous):

3/8=x/12 solve for x x=9/2

OpenStudy (anonymous):

and so on . . .

OpenStudy (anonymous):

2x/16x + 30 = -1/x x=-8/241